solve each equation for the specified variable.
T = C(2+AB) solve for B
so far this is what i have..
divide both sides by c.. C cancels out so your left with T=(2+AB)
Subtract both sides by 2.. 2 cancels out so your left with (T-2)/ (AB)
thats as far as i can understand as to what to do.

Answer by rfer(16322)

t=c(2+ab)
t/c=2+ab
t/c-2=ab
(t/c-2)/a=b